OpenStudy (christos):

Find the x -intercept and y -intercept of f(x) = (x - 4)^2(x+5)^2

hartnn (hartnn):

To get x-intercept put y=0 in the equation f(x) = (x - 4)^2(x+5)^2=y to get y-intercept put x=0 in the equation f(x) = (x - 4)^2(x+5)^2=y

OpenStudy (christos):

ok doing it atm

OpenStudy (christos):

first equation so far is 2x^3 +x^4 - 39x^2 - 40x + 400 = 0 Do I follow the correct path...?

hartnn (hartnn):

sorry to say no :P

hartnn (hartnn):

(x - 4)^2(x+5)^2=0 can't you directly say, x=4,4,5,5 like if (x-a)=0, x=a (x-a)(x-b)=0, then x=a,b (x-a)^2=0, then x=a,a

OpenStudy (christos):

x= -4 , 4, -5, 5 ??

hartnn (hartnn):

sorry, it'll be, (x - 4)^2(x+5)^2=0 x=4,4,-5,-5 why -4 and 5 ??

OpenStudy (christos):

I mean why did you type 4 and 5 2 times

hartnn (hartnn):

(x - 4)^2(x+5)^2=0 (x-4)^2 = 0 , (x-5)^2=0, (x-4)=0 , (x-4)=0 , (x-5) =0, (x-5)=0 ok, (x-a)^2=0 is a Quadratic Equation and it MUST have 2 roots. in this case, (x-a)^2=0, have 2 SAME(equal) roots as 'a' so we can say that its ha only 1 root 'a' but more precisely, it has 2 equal roots, a,a

hartnn (hartnn):

for the intercept case, x-intercept = 4,4 implies that the curve f(x) cuts the point (4,0) TWICE.

OpenStudy (christos):

I see

hartnn (hartnn):

doubts ?

hartnn (hartnn):

and whats y-intercept=... ?

OpenStudy (christos):

4 4 -5 -5 400 :P

hartnn (hartnn):

\(\huge \color{green}\checkmark\)